# Overcoming Long-Context Limitations of State-Space Models via Context-Dependent Sparse Attention

### v1

@@ 2.2 Locality-Sensitive Hashing Attention
[…]
@@ 4.1 Limitation of LSH Attention
Figure 3: The hybrid architectures of HAX integrated to Mamba (Gu & Dao, 2024) and Mamba2 (Dao & Gu, 2024). "SA" is short for "sparse attention", which is based on the sparse attention pattern defined in Eq. 21.
@@ 4.2 Key Selection (KS) Attention
With an ideal key scoring module that assigns the highest scores to the globally important keys, KS attention effectively covers $k$vertical-stripes" within attention patterns.
@@
where ${\mathbf{K}}[\mathcal{I}]$ is the selected key representations, ${\mathbf{M}}[\mathcal{I}]$ is the auto-regressive mask restricted to those positions, and $\sigma(\cdot)$ is the sigmoid function. With predicted scores $x\in\mathbb{R}^{k}$ and target scores $y\in\mathbb{R}^{k}$, we construct pairwise logits and targets:
@@ 5 Experiments
Table 1: Results of Ruler benchmark. We compare different sparse attention integrated with Mamba, including CISA methods: dilated attention (D), sliding window attention (SW), and their combination (SW+D), and A-shaped attention (A), as well as CDSA methods: LSH attention, and our proposed key selection (KS) attention and HAX. The best average performance is in bold.
@@
Mamba 99.6 53.4 7.4 20.8 0 0 17.8 4.8 1.5 1.3 100 14 26.7
+D 95.0 27.6 3.2 19.8 0 0 10.1 4.6 2.8 1.6 100 12.4 23.1
+SW 100 74 8 21.6 0.4 0 20.3 7.0 1.7 1.28 100 10.6 28.7
+SW+D 99.8 54.2 6.4 22 0 0 17.6 5.0 4.2 1.3 100 14 26.9
+A 99.8 63 4.8 24 0 0 19.8 3 2.2 1.4 100 12.2 27.5
+LSH 99.8 36.2 12.2 24 0 0 12.3 3.4 4.9 1.7 100 16.3 25.9
+KS (ours) 99.6 58.8 14.6 23.2 0 0 20.9 11 2.2 1.1 100 11.8 28.6
+HAX (ours) 100 92.4 34.6 24 0.2 0 20.4 3.8 4.9 1.7 100 12.8 32.9
@@
Mamba 6.11 15.16 3.35 34.57 16.73 12.72 2.51 3.02 0.85 0.5 4.97 16.49 35.76 1.67 10.5 12.76 11.10
+D 4.26 10.22 2.56 32.1 11.16 9.63 2.05 2.08 0.76 1.13 4.34 12.72 32.95 2.85 14.5 12.87 9.76
+SW 5.57 15.03 3.34 34.63 16.06 11.38 2.13 2.86 1.99 0.5 6.34 16.05 35.4 1.74 12 13.14 11.14
+SW+D 6.77 14.78 2.89 34.7 16.72 12.03 2.11 2.83 0.95 0.5 5.04 16.72 35.99 1.72 10.5 14.07 11.15
+A 6.58 15.12 3.26 34.77 16.33 12.26 2.2 2.96 1.69 0.42 5.47 15.58 35.5 1.75 10.5 13.55 11.12
+LSH 3.15 9.18 1.77 26.06 9.51 5.66 1.33 1.52 1.17 0.38 3.34 8.49 26.34 1.51 10.75 8.44 7.41
+KS (ours) 6.32 15.92 3.22 34.1 17.11 12.03 2.56 2.71 0.95 0 5.34 16.46 35.57 1.67 10.5 13.6 11.13
+HAX (ours) 6.71 15.47 3.34 34.81 14.95 13.64 1.76 2.6 1.64 0.3 5.54 14.22 35.47 1.89 14 15.17 11.34
@@ 5.1 Empirical Verification on Multi-Query Joint Recall
Data. We construct a multi-query joint recall dataset in which the number of context blocks and the number of keys per context are independently sampled from the range $[5,16]$, and the size of the vocabulary is fixed at $|\mathcal{V}|=16$. The dataset consists of 1M training examples, along with 10K samples each for validation and testing.
@@
[…]
Results. The results are summarized in Table 5. Compared to the Mamba and Mamba2 base architectures, most hybrid sparse attention models show performance improvements. In particular, our proposed HAX model consistently achieves the best performance, surpassing the base architectures by more than 100%. These findings underscore the effectiveness of our approach.
@@ 5.2 Continual Pre-training on Natural Language
Setup. To evaluate our method on real-world long-context natural language data, we perform continual pre-training based on the publicly released Mamba 130M checkpoint (Gu & Dao, 2024). As in section 5.1, we augment the Mamba architecture with sparse attention, as illustrated in Figure 3. We include SSMs integrated with CISA as baselines, along with ablated variants of our HAX model.
Validation Loss during Continual Pre-Training. Figure 5 shows the validation loss $\mathcal{L}_{LM}$ of continual pre-training. As observed, the Mamba base architecture and all baseline variants exhibit either training instability or plateau early in the training process. In contrast, our proposed HAX model is the only architecture that shows a consistent decline in validation loss throughout the training process, indicating improved stability and sustained learning.
Ruler and LongBench Evaluation. Ruler (Hsieh et al., 2024) is a synthetic long-context NLP benchmark designed to assess model performance on tasks including retrieval, multi-hop reasoning, aggregation, and question answering. LongBench (Bai et al., 2024) comprises real-world NLP long-context tasks, including question answering, summarization, few-shot learning, retrieval, aggregation, and code completion. We perform instruction tuning after the continual pre-training stage following Mamba-Chat(Mattern & Hohr, 2023), and then evaluate the models after instruction tuning. The evaluation results, summarized in Tables 1 and 2, show that among all hybrid sparse attention variants, our proposed HAX model is the only one that outperforms the Mamba baseline by a significant margin on average.
@@ Appendix A Relationship Between the Argmax and Sign-Bit LSH Binning Rules
[…]
=$\mathcal{B}=\left\{{\mathbf{B}}_{\mathbf{b}}={\textstyle\sum_{j=1}^{h}b_{j}{\mathbf{H}}_{j}}\Bigm{|}\mathbf{b}=(b_{1},\dots,b_{h})\in\{-1,+1\}^{h}\right\}\subset\mathbb{R}^{d}.$ (23)
@@ Appendix B Theoretical Proof
where ${\mathbf{c}}$ is the representation of the current associated context token, ${\mathbf{k}}$ is the representation of the current associated key token, and ${\mathbf{v}}$ is the representation of the nearest value token. $is\_v$ is a sign indicator ($-1$ or $1$) that specifies whether the current token is a value token.
@@
where ${\mathbf{c}}$ is the constructed representation of this context token on the unit sphere, and the final coordinate is set to $-1$ to indicate that the current token is not a value. Similarly, a key token is embedded as
@@
In the second layer, LSH attention operates on the SSM state $[{\mathbf{c}},{\mathbf{k}},{\mathbf{v}},is\_v]\in\mathcal{U}$, using $[{\mathbf{c}},{\mathbf{k}},{\mathbf{0}},is\_v]$ as the LSH attention key representation, $[{\mathbf{c}},{\mathbf{k}},{\mathbf{0}},1]$ as the LSH attention query representation, and $[{\mathbf{0}},{\mathbf{0}},{\mathbf{v}},1]$ as the LSH attention value representation. This design ensures that value tokens in the information component and key tokens in the inquiry component that share the same context and key (i.e., identical ${\mathbf{c}}$ and ${\mathbf{k}}$ representations in the SSM state) will always be assigned to the same hash bin. With a sufficient number of, e.g. $O(n)$ hash bins, which can be efficiently constructed using sign-bit binning rule with $O(\log n)$ random projection vectors, values associated with each key in every specified context are reliably retrievable by LSH attention. This step is the bottleneck of computation with $O(n\log^{2}n)$ time complexity. ∎
@@ C.2 Expressiveness of SSMs Integrated with CDSA on Multi-Level Context Joint Recall
where ${\mathbf{z}}_{i}$ denotes the representation of the nearest $i$-th level context token, ${\mathbf{v}}$ represents the nearest value token, and $is\_v$ is a sign indicator ($-1$ or $1$) that specifies if the current token is a value token.
@@
where ${\mathbf{z}}_{i}$ is the context token representation, and the final coordinate is set to $-1$ to indicate that the token is not a value. Similarly, a value token is embedded as
@@
In the second layer, LSH attention operates on the SSM state $[{\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w},{\mathbf{v}},is\_v]$, using $[{\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w},{\mathbf{0}},is\_v]$ as the LSH attention key representation, $[{\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w},{\mathbf{0}},1]$ as the LSH attention query representation, and $[{\mathbf{0}},{\mathbf{0}},...,{\mathbf{v}},1]$ as the LSH attention value representation. This design ensures that tokens that share the same context (i.e., identical $[{\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w}]$ in the SSM state) will always be assigned to the same hash bin. With a sufficient number of, e.g. $O(n)$ hash bins, which can be efficiently constructed using sign-bit binning rule with $O(\log n)$ random projection vectors, value representations associated with identical key combinations in the context are reliably retrievable by LSH attention. This step is the bottleneck of computation with $O(wn\log^{2}n)$ time complexity. ∎
@@ D.1 Empirical Verification on Joint Recall
For all models, we fix the number of layers to $2$, set the hidden size to $64$, and use $k=64$. For variants that integrate multiple sparse attention components, namely SW+D, A (consists of a SW component and a sink attention component, a sink attention always attends to the first $k$ tokens in the sequence only), and HAX (LSH+KS), we allocate $k=32$ to each component, in order to maintain a global $k=64$. For both LSH and HAX (LSH+KS), we adopt the sign-bit binning strategy (Eq. 8) with $h=8$, and refresh the random hashing matrix at each training step. To ensure a fair comparison, we double the hidden size of the Mamba and Mamba2 baselines, which do not include any sparse attention mechanism, to $128$. Additionally, we include a Samba baseline, consisting of $2$ Mamba layers and $2$ sliding window attention layers. For Samba, the hidden size and sliding window width $(k)$ are both set to $64$. We use AdamW optimizer. All models are trained for $400,000$ steps with a batch size of $64$. Our implementation is based on Flash-Linear-Attention (Yang & Zhang, 2024).
@@ D.2 Continual Pre-training on Natural Language
Following Mamba (Gu & Dao, 2024), we adopt the uncopyrighted the Pile dataset (Gao et al., 2021) as our pre-training corpus. To enhance the long-context modeling capability, we filter samples to retain only those with tokenized lengths of at least $4,096$ tokens. After continual pre-training based on the publicly released Mamba 130M checkpoint, we select the checkpoint with the lowest validation loss and perform instruction tuning on the UltraChat dataset (Ding et al., 2023b), following Mamba-Chat pipeline (Mattern & Hohr, 2023). Finally, we evaluate the instruction-tuned model on the Ruler (Hsieh et al., 2024) and LongBench (Bai et al., 2024) benchmarks.
For all experiments, we fix the sparsity parameter at $k=128$. For variants that integrate multiple sparse attention components, namely SW+D, A, and HAX (LSH+KS), we set $k=64$ to each component, in order to maintain a global $k=128$. For both LSH and HAX (LSH+KS), we adopt the argmax binning strategy (Eq. 7) with $h=k$. We resample the random hashing matrix at each training step and fix a random hashing matrix during evaluation. To visualize the validation loss, we perform continual pre-training for $50,000$ steps with a context length of $2,048$ and a global batch size of $64$. For downstream evaluation, we conduct continual pre-training for $10,000$ steps with a context length of $3,072$ and the same global batch size. This is followed by instruction tuning for $3,000$ steps, also with a global batch size of $64$. We use AdamW optimizer.
At the beginning of continual pre-training, the ${\mathbf{K}}$ and ${\mathbf{Q}}$ projection weights are initialized using the parameters of the ${\mathbf{B}}$ and ${\mathbf{C}}$ projections from Mamba, respectively, based on state-space duality(Dao & Gu, 2024). The gating parameters ${\mathbf{E}}$ are initialized to zero. During continual pretraining, all models follow the cosine learning rate schedule used in Mamba, with a maximum learning rate of $3$e-$4$ and a minimum of $1$e-$5$. A warm-up phase of $200$ steps with a learning rate of 00 precedes the cosine schedule. For instruction tuning, we apply a $200$-step zero learning rate phase followed by $800$ steps of linear warm-up, after which the learning rate remains constant. The peak learning rate during instruction tuning is set to $3$e-$6$ for Mamba+D, Mamba+LSH, and Mamba+HAX, and $1$e-$5$ for all other architectures.

### v2

@@ 2.2 Locality-Sensitive Hashing Attention
Next, a random projection matrix ${\mathbf{H}}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)\in\mathbb{R}^{d\times h}$ is sampled to project the normalized vectors $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$ into the hash space. Then, we consider two binning rules which assign each vector $\tilde{{\mathbf{Q}}}_{i}$ and $\tilde{{\mathbf{K}}}_{i}$ to a discrete hash bin: the argmax binning rule and the sign-bit binning rule.
@@ 4.1 Limitation of LSH Attention
Figure 3: Illustration of the HAX architecture, and Mamba (Gu & Dao, 2024) integrated with HAX.
@@ 4.2 Key Selection (KS) Attention
With an ideal key scoring module that assigns the highest scores to the globally important keys, KS attention effectively covers $k$ "vertical-stripes" within attention patterns.
@@
where $\mathbf{K}$$\mathcal{I}$] is the selected key representations, $\mathbf{M}$$\mathcal{I}$] is the auto-regressive mask restricted to those positions, and $\sigma(\cdot)$ is the sigmoid function. With predicted scores $x\in\mathbb{R}^{k}$ and target scores $y\in\mathbb{R}^{k}$, we construct pairwise logits and targets:
@@ 5 Experiments
Table 1: Results of Ruler benchmark at 2K context length. We compare different sparse attention integrated with Mamba, including CISA methods: dilated attention (D), sliding window attention (SW), and their combination (SW+D), and A-shaped attention (A), as well as CDSA methods: LSH attention, and our proposed key selection (KS) attention and HAX. The best average performance is in bold.
@@
Mamba 100 95.2 86.8 34.4 2.6 0 32.4 28.35 2.76 1.4 100 25.2 42.43
+D 0 0 0 0 0 0 0 0 17.4 0 100 0 9.78
+SW 100 93.6 83.6 36.2 3.4 0 31.05 31.45 5.24 1.26 100 24.4 42.52
+SW+D 100 95 82.4 34 2.6 0 32.9 29.45 4.72 1.22 100 24.8 42.26
+A 100 93.2 81.4 34.4 2.6 0 31.7 27.1 4.4 1.26 100 26.8 41.91
+LSH 0 0 0 0 0 0 0 0 0.2 0.06 100 0 8.36
+KS (ours) 100 94.4 84.4 33.6 1.8 1.8 37.25 33.85 4.04 1 100 25 43.10
+HAX (ours) 100 98.8 89.6 45.4 3.8 0.2 35.5 28.2 2.04 1.64 100 29.6 44.57
@@
Mamba 13.39 20.3 8.37 39.51 20.36 27.62 3.89 8.35 1.04 1.5 10.69 20.36 38.47 20.15 45 37.59 19.79
+D 0 0.21 0 2.87 0.26 0 0 0 0 0 0 0.72 3.46 0.7 0 0 0.51
+SW 13.81 20.58 9.25 38.06 20.64 27.55 5.15 8.12 1.2 1.5 10.55 19.57 38.38 22.42 43 36.55 19.77
+SW+D 13.6 20.39 8.55 39 20.74 27.55 4.33 7.98 0.83 1.56 10.17 19.73 38.84 22.49 43 37.02 19.74
+A 13.63 20.77 8.55 38.59 20.2 27.31 4.56 7.66 0.27 1.74 11.32 20.09 38.5 20.73 44.5 37.16 19.72
+LSH 1.66 1.49 1.5 7.74 0.98 2.32 0.96 0.87 0.1 0 2.3 5.05 7.13 3.89 0 0.9 2.31
+KS (ours) 12.93 20.93 9.34 38.22 20.83 26.56 4.34 7.46 0.27 1.67 11.65 19.35 38.17 22.97 45 37.52 19.83
+HAX (ours) 12.73 20.64 8.79 39.3 21.79 28.31 3.95 7.64 0.83 2 11.94 19.63 39.07 21.04 43 38.47 19.95
@@ 5.1 Empirical Verification on Multi-Query Joint Recall
Data. We construct a multi-query joint recall dataset in which the number of context blocks and the number of keys per context are independently sampled from the range $[5,16]$, and the size of the vocabulary is fixed at $|\mathcal{V}|=16$. The dataset consists of 1.4M training examples, along with 14.4K samples each for validation and testing.
@@
Setup. For evaluation, we calculate mean accuracy per-sample and average over the testset. For each base architecture and hybrid sparse attention variant, we conduct experiments using $3$ different random seeds for each learning rate in $\{3$e-33, 11e-33, 33e-4}4\}, and report the average performance corresponding to the best-performing learning rate.
Results. The results are summarized in Table 5. Compared to the Mamba and Mamba2 base architectures, most hybrid sparse attention models show performance improvements. In particular, our proposed HAX model consistently achieves the best performance, which underscores the effectiveness of our approach.
@@ 5.2 Continual Pre-training on Natural Language
Setup. To evaluate our method on real-world long-context natural language data, we perform continual pre-training based on the publicly released Mamba checkpoints (Gu & Dao, 2024). As in section 5.1, we augment the Mamba architecture with sparse attention, as illustrated in Figure 3. We include SSMs integrated with CISA as baselines, along with ablated variants of our HAX model. Overall, our HAX adds only about 1% additional parameters to the released Mamba architecture. All models are continually pre-trained for 50K steps with a context length of 2K.
Validation Loss during Continual Pre-Training. Figure 5 shows the validation loss $\mathcal{L}_{LM}$ of 130M model continual pre-training on The Pile (Gao et al., 2021). As observed, the Mamba base architecture and all baseline variants exhibit either training instability or plateau early in the training process. In contrast, our proposed HAX model is the only architecture that shows a consistent decline in validation loss throughout the training process, indicating improved stability and sustained learning.
Ruler and LongBench Evaluation. Tables 1 and 2 present the downstream performance for the Mamba 790M model after continual pre-training for 50K steps on TxT360 (Tang et al., 2024) followed by instruction fine-tuning for 10K steps on UltraChat (Ding et al., 2023b) following Mamba-Chat (Mattern & Hohr, 2023). Ruler (Hsieh et al., 2024) is a synthetic long-context NLP benchmark designed to assess model performance on tasks including retrieval, multi-hop reasoning, aggregation, and question answering. LongBench (Bai et al., 2024) comprises real-world NLP long-context tasks, including question answering, summarization, few-shot learning, retrieval, aggregation, and code completion. The evaluation results, summarized in Tables 1 and 2, show that among all hybrid sparse attention variants, our proposed HAX model is the only one that outperforms the Mamba baseline by a significant margin on average.
@@ Appendix A Relationship Between the Argmax and Sign-Bit LSH Binning Rules
Expanding the projection matrix. Let the original random projection be ${\mathbf{H}}=[{\mathbf{H}}_{1},...,{\mathbf{H}}_{h}]\in\mathbb{R}^{d\times h},{\mathbf{H}}_{j}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1).$ Define a codebook of $2^{h}$ signed prototypes
$\mathcal{B}=\left\{{\mathbf{B}}_{\mathbf{b}}={\textstyle\sum_{j=1}^{h}b_{j}{\mathbf{H}}_{j}}\Bigm{|}\mathbf{b}=(b_{1},\dots,b_{h})\in\{-1,+1\}^{h}\right\}\subset\mathbb{R}^{d}.$ (23)
@@ Appendix B Theoretical Proof
where $\mathbf{c}$} is the representation of the current associated context token, $\mathbf{k}$} is the representation of the current associated key token, and $\mathbf{v}$} is the representation of the nearest value token. is_vis\_v is a sign indicator (1-1 or $1$) that specifies whether the current token is a value token.
@@
where $\mathbf{c}$} is the constructed representation of this context token on the unit sphere, and the final coordinate is set to 1-1 to indicate that the current token is not a value. Similarly, a key token is embedded as
@@
In the second layer, LSH attention operates on the SSM state ${\mathbf{c}},{\mathbf{k}},{\mathbf{v}},is\_v]\in\mathcal{U}$, using ${\mathbf{c}},{\mathbf{k}},{\mathbf{0}},is\_v]$ as the LSH attention key representation, ${\mathbf{c}},{\mathbf{k}},{\mathbf{0}},1]$ as the LSH attention query representation, and ${\mathbf{0}},{\mathbf{0}},{\mathbf{v}},1]$ as the LSH attention value representation. This design ensures that value tokens in the information component and key tokens in the inquiry component that share the same context and key (i.e., identical $\mathbf{c}$} and $\mathbf{k}$} representations in the SSM state) will always be assigned to the same hash bin. With a sufficient number of, e.g. $O(n)$ hash bins, which can be efficiently constructed using sign-bit binning rule with $O(\log n)$ random projection vectors, values associated with each key in every specified context are reliably retrievable by LSH attention. This step is the bottleneck of computation with $O(n\log^{2}n)$ time complexity. ∎
@@ C.2 Expressiveness of SSMs Integrated with CDSA on Multi-Level Context Joint Recall
where ${\mathbf{z}}_{i}$ denotes the representation of the nearest $i$-th level context token, $\mathbf{v}$} represents the nearest value token, and is_vis\_v is a sign indicator (1-1 or $1$) that specifies if the current token is a value token.
@@
where ${\mathbf{z}}_{i}$ is the context token representation, and the final coordinate is set to 1-1 to indicate that the token is not a value. Similarly, a value token is embedded as
@@
In the second layer, LSH attention operates on the SSM state ${\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w},{\mathbf{v}},is\_v]$, using ${\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w},{\mathbf{0}},is\_v]$ as the LSH attention key representation, ${\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w},{\mathbf{0}},1]$ as the LSH attention query representation, and ${\mathbf{0}},{\mathbf{0}},...,{\mathbf{v}},1]$ as the LSH attention value representation. This design ensures that tokens that share the same context (i.e., identical ${\mathbf{z}}_{1},{\mathbf{z}}_{2},...,{\mathbf{z}}_{w}]$ in the SSM state) will always be assigned to the same hash bin. With a sufficient number of, e.g. $O(n)$ hash bins, which can be efficiently constructed using sign-bit binning rule with $O(\log n)$ random projection vectors, value representations associated with identical key combinations in the context are reliably retrievable by LSH attention. This step is the bottleneck of computation with $O(wn\log^{2}n)$ time complexity. ∎
@@
Table 3: Results of Ruler benchmark at 4K context length. We compare different sparse attention integrated with Mamba, including CISA and CDSA methods as in Tab. 1. The best average performance is in bold.
Model NIAHS1 NIAHS2 NIAHS3 NIAHMK1 NIAHMK2 NIAHMK3 NIAHMV NIAHMQ VT CWE FWE QA1 QA2 Average
Mamba 100 19.2 9.2 5.2 0.6 0 0.65 1.85 5.72 4.3 100 26.2 78.6 27.04
+D 0 0 0 0 0 0 0 0 65.2 0 100 0 73 18.32
+SW 100 18.2 5 5.6 0.2 0.2 1.75 2.1 4.4 4.78 100 22.4 78.8 26.42
+SW+D 100 20.2 8.2 6.2 0.4 0.2 1.45 1.25 4.36 4.32 100 24.6 79 26.94
+A 100 19.8 7.2 6.6 1.2 0.2 1.3 0.8 4.36 3.24 100 24.2 79 26.76
+LSH 0 0 0 0 0 0 0 0 0 0.04 100 0.2 73 13.33
+KS (ours) 99.8 20.2 7.8 8.2 1 0.2 1.2 1.85 9.32 2.86 100 22.6 78.6 27.20
+HAX (ours) 100 23 11 9 0.8 0 3.4 2.75 6.36 2.54 100 29 78.8 28.20
@@ D.1 Empirical Verification on Joint Recall
For all models, we fix the number of layers to 2, set the hidden size to 64, and use $k=64$. For variants that integrate multiple sparse attention components, namely SW+D, A (consists of a SW component and a sink attention component, a sink attention always attends to the first $k$ tokens in the sequence only), and HAX (LSH+KS), we allocate $k=32$ to each component, in order to maintain a global $k=64$. For both LSH and HAX (LSH+KS), we adopt the sign-bit binning strategy (Eq. 8) with $h=8$, and refresh the random hashing matrix at each training step. To account for the absence of sparse attention mechanisms, we additionally evaluate Mamba and Mamba2 baselines with their hidden size doubled to 128, ensuring a fair comparison. Additionally, we include a Samba baseline, consisting of 2 Mamba layers and 2 sliding window attention layers. For Samba, the hidden size and sliding window width $(k)$ are both set to 64. We use AdamW optimizer. All models are trained for 400,000 steps with a batch size of 64. Our implementation is based on Flash-Linear-Attention (Yang & Zhang, 2024).
@@ D.2 Continual Pre-training on Natural Language
For all experiments, we fix the sparsity parameter at $k=128$. For variants that integrate multiple sparse attention components, namely SW+D, A, and HAX (LSH+KS), we set $k=64$ to each component, in order to maintain a global $k=128$. For both LSH and HAX (LSH+KS), we adopt the argmax binning strategy (Eq. 7) with $h=k$. We resample the random hashing matrix at each training step and fix a random hashing matrix during evaluation.
To enhance the long-context modeling capability, we filter samples to retain only those with tokenized lengths of at least 2,048 tokens. At the beginning of continual pre-training, the $\mathbf{K}$} and $\mathbf{Q}$} projection weights are initialized using the parameters of the $\mathbf{B}$} and $\mathbf{C}$} projections from Mamba, respectively, based on state-space duality (Dao & Gu, 2024). During continual pretraining, we use a cosine schedule with a maximum learning rate of 3e-4 for 130M models and 1.5e-4 for 790M models, and a minimum learning rate of 1e-5. A warm-up phase of 200 steps with a learning rate of 0 followed by 800 steps of linear warm-up precedes the cosine schedule. For instruction tuning, we also apply a 200-step 0 learning rate phase followed by 800 steps of linear warm-up, after which the learning rate remains constant at 3e-6. For both continual pre-training and instruction-tuning, we use AdamW optimizer and train with a global batch size of 64 and a context length of 2K.